# Impacts of Dynamical Dark Energy on the Neutrino Mass Constraints

### v1

@@ Abstract
The difference between the total neutrino mass estimates ($\sum m_{\nu}$) derived from cosmological data within the standard $\Lambda$CDM model and those obtained from terrestrial particle physics experiments underscores the need to explore alternative scenarios. Recent analyses have shown that a dynamic dark energy modeled by the CPL parameterization of the dark energy equation of state (EoS) can ease the constraints on $\sum m_{\nu}$, thus alleviating this tension. This study investigates the robustness of this discrepancy by assessing the extent to which the CPL assumption influences the results. We examine how other EoS parameterizations, such as the Barboza-Alcaniz (BA) and Jassal-Bagla-Padmanabhan (JBP) parameterizations, affect $\sum m_{\nu}$ estimates. We perform a Markov Chain Monte Carlo (MCMC) analysis combining the latest baryon acoustic oscillation data from DESI with the Planck 2018 cosmic microwave background data, which includes information on temperature, polarization, and lensing, as well as the Pantheon+ Type Ia supernovae observations. While both the BA and JBP parameterizations can also resolve the tension, our results show a correlation between the dark energy EoS and the constraints on neutrino mass.
@@ I Introduction
In this context, various alternative proposals have been suggested to explain the current cosmic acceleration beyond the cosmological constant (see e.g. [5, 6, 7, 8, 9, 10, 11]). In general, such proposals assume the existence of scalar fields that have not yet reached their ground state or modifications in the Friedmann equations, implying time-dependent EoS parameters $w(a)$. Phenomenologically, several authors have also proposed $w(a)$ parameterization [12, 13, 14, 15, 16], which insert the $\Lambda$CDM model in a more general framework, opening the possibility of checking deviations from the standard case.
To assess the observational viability of dynamical dark energy (DDE) models, several probes have been utilized from both the early and local Universe [17, 18, 19, 20, 21, 22, 23]. It is important to note that CMB data alone cannot constrain such models tightly. This limitation is due to a geometric degeneracy between the parameters that affect the late-time Cosmology and the angular diameter distance to the surface of the last scattering. Therefore, this degeneracy must be broken using local Universe observables to constrain this class of models. Although incorporating information such as SN data alongside CMB observations helps tighten the constraints, no preference for this class of models has been shown at first. This remained the case until very recently, when BAO observations from the Dark Energy Spectroscopic Instrument (DESI), combined with Planck CMB and Pantheon+ SN data, showed a significant preference for DDE using the Chevallier-Polarski-Linder (CPL) EoS parameterization, $w(a)=w_{0}+w_{a}(1-a)$ [24]. Even more recently, it has been shown that this preference is not limited to the CPL but extends to other parameterizations as well. This result further reinforces the preference indicated by parametric analyses of the DESI and SNe data for DDE models [25] [^1] This preference for a DDE is not confirmed by non-parametric (model-independent) analyses of the current BAO and SN data, as shown in [26, 27] (see also [28])..
@@
Back to the cosmological perspective, neutrinos have a well-marked role in the expansion of the cosmos. They act as radiation in the very early universe, later changing to a hot dark matter behavior after the non-relativistic transition. At the perturbative level, the neutrinos dispersion acts to suppress the growth of perturbation at small scales, characterized by their free-streaming and, ultimately, by the sum of their masses. One of the current tightest constraints, $\sum m_{\nu}<0.07\;\text{eV}$ at 95% confidence level (C.L), comes from the combination of CMB and DESI BAO data within the $\Lambda$CDM framework, showing a clear preference for the NO [24, 30] (see also [31]). More recently, an upper limit of $\sum m_{\nu}<0.042$ eV was obtained using CMB data combined with DESI BAO, the galaxy cluster angular diameter distance, and an $H_{0}$ prior from SH0ES [32].
These constraints are significantly relaxed within a DDE framework modeled by the CPL parameterization, yielding $\sum m_{\nu}<0.15\;\text{eV}$ at 95% C.L [24]. Therefore, the current cosmological limits for the $\Lambda$CDM are on the verge of excluding both inverted and normal ordering, thereby generating tension with the results from particle physics experiments [33, 32]. This work examines how other well-established EoS parameterizations, such as the Barboza-Alcaniz (BA) and Jassal-Bagla-Padmanabhan (JBP) parameterizations, impact the $\sum m_{\nu}$. In our analysis, we perform a Markov Chain Monte Carlo (MCMC) analysis combining the latest CMB, SN, and DESI BAO data. The paper is organized as follows. In Section 2, we detail the functional forms of the DDE models considered. In Section 3, we present the methodology adopted and the results obtained. In Section 4, we present and discuss our main conclusions.
@@ II Dynamical Dark Energy Models
The dark energy also affects the adiabatic sound speed of the fluid $c_{s}^{2}$. To investigate the impact of of DDE models on the sum of neutrino mass, we will use two functional forms for the EoS parameter:
@@
    $w(a)=w_{0}+w_{a}\frac{1-a}{a^{2}+(1-a)^{2}}$ (5)
This parameterization exhibits a linear behavior at low-redshifts. Unlike the CPL parameterization, which blows up exponentially as the scale factor approaches infinity (for $w_{a}>0$), the BA parameterization remains well-behaved at all times, while still allowing for deviations from the CPL scenario [^2] Among other well-known EoS parameterizations, the BA gives the most significant improvement over the $\Lambda$CDM model when analyzed in the context of current CMB, BAO, and SN observations [25]..
@@ III Analysis and Results
To obtain constraints on the cosmological parameters for the DDE models above, we perform a Monte Carlo Markov Chain (MCMC) numerical analysis using the Planck 2018 temperature and polarization data (including low and high-multipole modes and lensing) [34], along with SN data from Pantheon+ [35, 36] and geometric BAO information from the DESI collaboration [24]. For this, we implement the two parameterizations in the Boltzmann solver code Cosmic Linear Anisotropy Solving System (CLASS) [37], in which we interfaced with the Cobaya code to perform the MCMC statistical analysis [38, 39].
Figure 1: Confidence contours at $68\%$ and $95\%$ C.L. for the JBP and BA parameterizations with Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI BAO data.
Figure 1 shows the 68% and 95% confidence level contours obtained from our analysis. It is immediately evident that the main differences between them lie in the EoS parameters. It shows significant differences, particularly in $w_{0}$, $w_{a}$, and $\sum m_{\nu}$. Table 1 presents each model’s mean values and associated errors. The $2\sigma$ limits are provided for the sum of the neutrino masses, while the $1\sigma$ limits are shown for the other parameters. For the BA model, we find an upper 95% one-tail limit of $\sum m_{\nu}<0.18$ eV. The JBP model provides a more restrictive upper limit, with $\sum m_{\nu}<0.13$ eV. Using the same data set, the CPL parameterization yields an upper bound of $\sum m_{\nu}<0.18$ eV  [^3] Our neutrino mass constraints for the CPL model are slightly different from the one reported in the DESI paper [24]. Here, we use CMB data from the Planck baseline likelihood only, while the DESI collaboration uses Planck combined with Data Release 6 of the Atacama Cosmology Telescope.. These results demonstrate that the $\sum m_{\nu}$ constraints depend on the parameterization adopted in the analysis. However, for those considered in this study, the limits obtained alleviate the tension and agree with the mass hierarchy scheme [^4] For completness, we also conduct analyses using a logarithmic parameterization defined as $w(a)=w_{0}+w_{a}\ln(a)$. Although not presented here, the results derived from this approach are similar to those obtained by the BA parameterization..
   Parameter BA JBP CPL
$h$ $0.6808\pm 0.0073$ $0.6792\pm 0.0071$ $0.6801\pm 0.0071$
$\Omega_{m}$ $0.3080\pm 0.0071$ $0.3082\pm 0.0069$ $0.3085\pm 0.0069$
$\Omega_{b}$ $0.0484\pm 0.0011$ $0.0486\pm 0.0010$ $0.0484\pm 0.0010$
$n_{s}$ $0.9657\pm 0.0040$ $0.9667\pm 0.0038$ $0.9658\pm 0.0037$
$\sigma_{8}$ $0.816\pm 0.012$ $0.814\pm 0.011$ $0.815^{+0.012}_{-0.011}$
$\sum m_{\nu}$$(\mathrm{eV})$ $0.072^{+0.11}_{-0.074}$$(2\sigma)$ $0.050^{+0.083}_{-0.051}$$(2\sigma)$ $0.070^{+0.11}_{-0.073}$$(2\sigma)$
$w_{0}$ $-0.858\pm 0.055$ $-0.795\pm 0.089$ $-0.813\pm 0.064$
$w_{a}$ $-0.36^{+0.17}_{-0.13}$ $-1.30\pm 0.60$ $-0.73^{+0.34}_{-0.26}$
@@
In Figure 2, we show the $w_{0}$-$w_{a}$ plane. When treating neutrino mass as a free parameter, the JBP model progressively aligns with the $\Lambda$CDM scenario, in contrast to the BA model, which remains more than $2\sigma$ away. Similarly, the CPL model shows that allowing for free neutrinos has a slight influence on its analysis, consistent with the results reported by the DESI collaboration. This result suggests that relaxing constraints on the dark energy parameters may help constrain the neutrino mass, although the correlation is not particularly strong.
Figure 3 shows the EoS evolution with redshift for each model, using Planck CMB temperature, polarization, and lensing data, along with Pantheon+ SN and DESI BAO. They behave similarly at low redshifts, remaining in the quintessence region, i.e., $w(z)>-1$. At high redshifts, there are significant differences. The BA model crosses at a slightly lower redshift and remains in the phantom region near $w(z)=-1$. Interestingly, for the JBP model, it tends to cross $w(z)=-1$ twice. It leaves the quintessence region and remains very close to $w(z)=-1$ with a trend toward the phantom region until it returns to the quintessential region for high $z$ values. These EoS behaviors for these models are similar to those presented in the work [25], where a similar analysis was performed with the neutrino mass fixed. The most significant difference is in the JBP behavior, which shows a larger error in $w(z)$.
Figure 2: $w_{0}-w_{a}$ plane for the JBP, BA and CPL parameterizations with Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI BAO data. The dashed line intersection represents the SCM values, $w_{a}=0$ and $w_{0}=-1$. The filled contours are for the models assuming the neutrino mass as a free parameter (Blue, Red and Green), while the black empty contours (shown here for completeness) stand for the JBP model with a fixed neutrino mass.
Figure 3: EoS evolution obtained for Planck+DESI+Pantheon+ data assuming the BA (Blue) and JBP parameterizations (red). The dashed line indicates the $\Lambda$CDM model, $w(z)=-1$.
@@ IV Final Remarks
Neutrinos are one of the most promising probes to unveil aspects of the physics at high energies. For instance, in the well-known seesaw paradigm [40, 41, 42, 43], the magnitude of the neutrino masses is dictated by a new energy scale, associated with lepton number violation. The mechanisms that lead to this symmetry breaking process may be associated with Grand Unification theories [44], and also may offer a solution to the fermion/anti-fermion unbalance of the universe [45]. Meanwhile, the micro-physical aspects of the standard (active) and sterile neutrinos are set by the Yukawa structure of the model’s Lagrangian, which may be constrained by the oscillation process observed in the atmospheric and solar neutrinos and also set the possible signatures left by new particles in related experiments [46, 47].
In this regard, the complementarity of cosmological and terrestrial observations is an important aspect for constraining the theory. While neither type of experiment can determine their fundamental properties independently, their combination can help narrow down the possible mass scale and provide insights into their interactions. Thus, as the tension between neutrino mass measurements derived from cosmological data in the $\Lambda$CDM framework and those obtained from precise ground-based experiments becomes increasingly apparent, it is crucial to explore alternative models to address this inconsistency. In this context, the DDE scenario modeled by the CPL parameterization has demonstrated the potential to resolve this discrepancy, yielding $\sum m_{\nu}<0.15\;\text{eV}$ at a 95% (CL) [24]. Additionally, parametric analyses of recent data from DESI indicate a slight preference for this model compared to the $\Lambda$CDM framework.
In this work, we investigated the robustness of the $m_{\nu}$ tension and the results obtained from the CPL model. By performing a MCMC analysis with the most recent available BAO data from DESI along with temperature, polarization, and lensing data from Planck and Type Ia SN from Pantheon+, we studied how different well-known EoS parameterizations, such as BA and JBP parameterizations, influence the estimates of $\sum m_{\nu}$. We found that the BA and JBP follow the same trend as the CPL by helping to relax the constraint on the sum of neutrino masses. The neutrino mass upper limit obtained for the BA is $\sum m_{\nu}<0.18$ eV while the JBP parameterization provides a more restrictive upper limit of $\sum m_{\nu}<0.13$ eV, a difference of $\sim 30\%$. We also note that the main difference lies primarily in the constraints on the EoS parameters. This result suggests that a larger uncertainty on $w_{0}$ and $w_{a}$ values leads to tighter constraints on the neutrino mass. For the JBP parameterization, we obtain constraints on $w_{0}$ and $w_{a}$ that are compatible with the $\Lambda$CDM model values within $2\sigma$. However, this is not the case for the BA, which remains more than $2\sigma$ away.
Finally, our results confirm the ongoing tension in the measurements of $\sum m_{\nu}$ within the context of the standard cosmological model, as reported by [33, 32]. They also demonstrate that dynamical dark energy models can effectively alleviate this issue. Furthermore, our analysis indicates that the degree to which constraints on $\sum m_{\nu}$ are relaxed depends moderately on the chosen parameterization.

### v2

# Thawing Dark Energy and Massive Neutrinos in Light of DESI

@@ Abstract
Recent analyses have shown that a dynamic dark energy modeled by the CPL parameterization of the dark energy equation of state (EoS) can ease constraints on the total neutrino mass compared to the standard $\Lambda$CDM model. This helps reconcile cosmological and particle physics measurements of $\sum m_{\nu}$. In this study, we investigate the robustness of this effect by assessing the extent to which the CPL assumption influences the results. We examine how alternative EoS parameterizations - such as Barboza-Alcaniz (BA), Jassal-Bagla-Padmanabhan (JBP), and a physically motivated thawing parameterization that reproduces the behavior of various scalar field models - affect estimates of $\sum m_{\nu}$. Although both the BA and JBP parameterizations relax the constraints similarly to the CPL model, the JBP parameterization still excludes the inverted neutrino mass hierarchy at $\sim 2.1\;\sigma$ with $\sum m_{\nu}<0.096$ eV. The thawing parameterization excludes the inverted hierarchy at $\sim 3.3\sigma$ and yields tighter constraints, comparable to those of the $\Lambda$CDM model, with $\sum m_{\nu}<0.071$ eV. Our analysis also reveals that the preference for unphysical negative neutrino masses is significantly reduced in all models considered. Finally, we show that the thawing model can be mapped into the BA and JBP $w_{0}$$w_{a}$ parameter space, with the apparent preference for the phantom regime actually supporting quintessence (non-phantom) models.
@@ I Introduction
In the constext of General Relativity, two major frameworks have emerged as alternative to model the Universe’s accelerated expansion (see e.g. [5, 6, 7, 8, 9, 10, 11]): the scalar field approach, exemplified by quintessence models, and phenomenological parameterizations of the dark energy equation of state (EoS), such as the widely used Chevallier–Polarski–Linder (CPL), Barboza–Alcaniz (BA), and Jassal–Bagla–Padmanabhan (JBP) forms [12, 13, 14, 15, 16]. In both frameworks, the dark energy EoS is a dynamical quantity, $w(z)$.
Quintessence models describe dark energy as a slowly rolling scalar field with a canonical kinetic term and a potential $V(\phi)$ that drives its evolution. A key feature of these models is that the EoS remains strictly above the phantom divide, $w(z)>-1$, unless the theory is extended to include exotic components. On the other hand, phenomenological parameterizations do not assume an underlying physical origin for DE but instead provide flexible, model-independent forms for $w(z)$. These can allow the EoS to cross the phantom divide and are useful for fitting cosmological data. However, as discussed in [17, 18, 19, 20, 21], a phantom-like behavior in parameterized EoS models — such as $w(z)<-1$ in the CPL framework — does not necessarily imply the existence of or even preference for phantom dark energy. When mapping physically motivated quintessence models into the CPL $w0$-$wa$ parameter space, many fall within the so-called phantom region, indicating that such behavior can emerge from standard scalar field dynamics. Therefore, while these parameterizations should not be dismissed — as they are useful in searching for deviations from $\Lambda$ — they are limited in their ability to reveal the physical nature of DE.
To evaluate the observational viability of dynamical dark energy (DDE) models, various probes from both the early and local Universe have been employed [22, 23, 24, 25, 26, 27, 28]. CMB data alone cannot strongly constrain these models due to a geometric degeneracy between late-time parameters and the angular diameter distance to the last scattering surface. Breaking this degeneracy requires the inclusion of local Universe observables. Although combining CMB with SN data improves constraints, no clear preference for DDE emerged initially. This remained the case until very recently, when BAO observations from the Dark Energy Spectroscopic Instrument (DESI), combined with Planck CMB and Pantheon+ SN data, showed a significant preference for DDE using the CPL EoS parameterization, $w(a)=w_{0}+w_{a}(1-a)$ [29]. Even more recently, it has been shown that this preference is not limited to the CPL but extends to other parameterizations as well. This result further reinforces the preference indicated by parametric analyses of the DESI and SNe data for DDE models [30].
@@
Back to the cosmological perspective, neutrinos have a well-marked role in the expansion of the cosmos. They act as radiation in the very early universe, later changing to a hot dark matter behavior after the non-relativistic transition. At the perturbative level, the neutrinos dispersion acts to suppress the growth of perturbation at small scales, characterized by their free-streaming and, ultimately, by the sum of their masses. One of the current tightest constraints, $\sum m_{\nu}<0.07\;\text{eV}$ at 95% confidence level (C.L), comes from the combination of CMB and DESI BAO data within the $\Lambda$CDM framework, showing a clear preference for the NO [29, 32] (see also [33]). More recently, the upper limit $\sum m_{\nu}<0.042$ eV was obtained using CMB combined with DESI BAO, galaxy cluster measurements, and a $H_{0}$ prior from SH0ES [34]. These bounds are substantially relaxed under a DDE framework modeled by CPL, yielding $\sum m_{\nu}<0.11\;\text{eV}$ at 95% C.L [29](see also [35]). Therefore, the current cosmological limits for the $\Lambda$CDM are on the verge of excluding both inverted and normal ordering, creating tension with the results of particle physics experiments [36, 34].
This work examines how other well-established EoS phenomenological parameterizations, such as Barboza-Alcaniz (BA) and Jassal-Bagla-Padmanabhan (JBP), affect the $\sum m_{\nu}$ constraints. We also consider a $w(z)$ parameterization that describes the overall behavior of several thawing quintessence models [21]. In our work, we performed a Markov Chain Monte Carlo (MCMC) analysis combining the latest CMB, SN, and DESI BAO data. We also perform a Bayesian analysis to compare the results of the above parameterizations with respect to the CPL parameterization. In Section 2, we detail the functional forms and the motivation of the DDE models considered. Section 3 presents the methodology adopted in our analysis and the results obtained. In Section 4, we present and discuss our main conclusions.
@@ II Dynamical Dark Energy Models
The dark energy also affects the adiabatic sound speed of the fluid $c_{s}^{2}$. To investigate the impact of DDE models on the sum of neutrino mass, we will use the following functional forms for the EoS,
@@
    $w(a)=w_{0}+w_{a}\frac{1-a}{a^{2}+(1-a)^{2}}\;.$ (5)
This parameterization exhibits a linear behavior at low-redshifts. Unlike the CPL parameterization, which blows up exponentially as the scale factor approaches infinity (for $w_{a}>0$), the BA parameterization remains well-behaved at all times, while still allowing for deviations from the CPL scenario.
  • The thawing quintessence parameterization [11, 21]
    $w(a)=-1+\alpha a^{\beta}\;.$ (6)
Standard quintessence scenarios generally predict an EoS parameter satisfying $w>-1$, never crossing into the phantom regime. To generalize a class of quintessence potentials, [37] proposed the ansatz, ${1}/{\rho_{\phi}}({d\rho_{\phi}}/{da})=-{3\alpha}/{a^{1+\beta}}$, which results in (6) when combined with the continuity equation, $\dot{\rho}_{\phi}+3H\rho_{\phi}[1+w(a)]=0$. For $\alpha>0$, Eq. (6) captures both thawing and freezing behaviors of scalar field models, with $\alpha$ setting the present-day EoS value and $\beta>0$ controlling the rate of deviation from $w=-1$. When both $\alpha>0$ and $\beta>0$, the model reproduces the typical dynamics of thawing quintessence, where the field is initially frozen due to Hubble friction and behaves like a cosmological constant, later evolving to $w(z)>-1$. For a recent analysis of this generalized quintessence scenario in the absence of neutrinos, see [21].
@@ III Analysis and Results
To obtain constraints on the cosmological parameters for the DDE models above, we perform a Monte Carlo Markov Chain (MCMC) numerical analysis using Planck 2018 temperature and polarization data, including low- and high-multipole measurements of temperature and polarization, combined with lensing (plikHM+TTTEEE+lowl+lowE+lensing) [38], along with Pantheon+ catalog SN data [39, 40] and geometric BAO information from the DESI collaboration data release two [29]. For this, we implement the parameterizations in the Boltzmann solver code Cosmic Linear Anisotropy Solving System (CLASS) [41], in which we interfaced with the Cobaya code to perform the MCMC statistical analysis [42, 43]. For the purpose of constraining the neutrino mass, we assume the standard degenerate mass spectrum, in which the three massive neutrino species each carry one third of the total mass, i.e., $\sum m_{\nu}=3(m_{\nu,1}=m_{\nu,2}=m_{\nu,3})$. See [44] for a detailed discussion on the choice of neutrino mass hierarchy in cosmological analyses. In our case, we also imposed a prior that the total neutrino mass must satisfy $\sum m_{\nu}>0$.
Figure 1: Constraints on $\sum m_{\nu}$ from Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI DR2 data assuming JBP, BA, and Thawing parameterizations. The dashed vertical line represents the inverted hierarchy $\sum m_{\nu}>0.10$ eV. For comparison, we also show the CPL prediction (black line).
Figure 1 shows the marginalized posterior distribution for $\sum m_{\nu}$ across the models explored in this work. For comparison, see Fig. 7 of [45], which presents results from the DESI collaboration using DESI DR2 BAO + CMB + Pantheon+ data. In their analysis of the CPL model, the posterior peaks at the prior boundary $\sum m_{\nu}=0$, suggesting that the data favor non-physical values of $\sum m_{\nu}<0$. This is not the case in our analysis: all models show a peak at $\sum m_{\nu}>0$, with only a small portion of the posterior excluded due to prior choice $\sum m_{\nu}>0$. It is important to note that DESI uses Planck combined with Data Release 6 of the Atacama Cosmology Telescope (ACT), while in our analysis we use CMB data from the Planck baseline likelihood only.
@@
[…]
Model $\sum m_{\nu}$ ($2\sigma$) $\Delta\ln{B}$
CPL $0.048^{+0.076}_{-0.049}$ eV $0.0$
BA $0.050^{+0.079}_{-0.051}$ eV $-0.64$
JBP $0.036^{+0.061}_{-0.037}$ eV $-0.60$
TWA $0.027^{+0.045}_{-0.027}$ eV $2.66$
Table 1: $2\sigma$ constraints on the sum of neutrino masses $\sum m_{\nu}$ and the Bayes factor, $\Delta\ln{B}$, for all models discussed in the text. Negative values of $\Delta\ln{B}$ favors the reference model (CPL).
In Figure 2, we show the $w_{0}$$w_{a}$ planes for JBP (upper panel) and BA parameterizations (bottom panel). When treating neutrino mass as a free parameter, the JBP model is consistent with the $\Lambda$CDM scenario within $2\sigma$, while the BA model lies nearly outside the $2\sigma$ boundary. Following the procedure of [20], we also project the $(\alpha,\beta)$ values from the thawing model into the $(w_{0},w_{a})$ parameter space by matching the expansion history $H(z)$. The shaded regions in both panels of Fig. 2 represent the $\alpha$$\beta$ values within the full prior range considered in the analysis: $\alpha\in[0,1]$ and $\beta\in[0,15]$. This shows that the phantom region in the $w_{0}$$w_{a}$ plane for both models is fully compatible with the scenario of thawing quintessence. The star markers correspond to the best-fit values, $\alpha=0.15$ and $\beta=7.84$, projected into the $w_{0}$$w_{a}$ parameter spaces for both models. These best-fit projections lie within the $1\sigma$ and $2\sigma$ confidence regions of the JBP and BA, respectively.
Figure 3 shows the EoS evolution with redshift for each model, using Planck CMB temperature, polarization, and lensing data, along with Pantheon+ SN and DESI BAO. The BA and JBP behave similarly at low redshifts, remaining in the quintessence region, i.e., $w(z)>-1$. The BA model crosses to phantom $w(z)<-1$ at a slightly lower redshift than the JBP, and remains in the phantom region near $w(z)=-1$. Interestingly, for the JBP model, it tends to cross $w(z)=-1$ twice, leaving the quintessence region and remaining very close to $w(z)=-1$ with a trend toward the phantom region until it returns to the quintessential region for high $z$ values. The thawing EoS remains exactly equal to $w(z)=-1$ at high redshifts and begins to evolve toward the $w(z)>-1$ region near $z\approx 0$.
Figure 2: $w_{0}-w_{a}$ plane for the JBP (upper) and BA (bottom) parameterizations obtained from Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI DR2 BAO data. The dashed line intersection represents the point, $w_{a}=0$ and $w_{0}=-1$. The shaded red (JBP) and blue (BA) regions represent the full range of $\alpha$$\beta$ priors projected into the $w_{0}$$w_{a}$ space, while the star markers indicate the projected best-fit $\alpha$$\beta$ values for each model. The red and blue dashed lines shows the dividing line between quintessence and phantom behavior for each EoS parameterizations.
Figure 3: EoS evolution obtained for Planck+DESI+Pantheon+ data assuming the BA (blue), JBP (red) and Thawing (green) parameterizations.
@@ IV Final Remarks
Our findings in this work highlight several key aspects. First, unlike previous results reported by the DESI collaboration [48], where the posterior for $\sum m_{\nu}$ in the CPL model peaks at the prior edge of $\sum m_{\nu}=0$, which indicates a preference for non-physical negative neutrino mass values, all models in our analysis peak at positive values. This suggests that Planck data when combined with ACT (used in DESI’s analysis) may significantly shift the results toward this negative preference. In addition, the thawing parameterization provides the strongest constraint on the neutrino mass, $\sum m_{\nu}<0.071$ eV at 95% C.L., comparable to that of $\Lambda$CDM and sufficient to exclude the inverted hierarchy at more than $3\sigma$. The JBP model also shows a mild preference for the normal hierarchy over the inverted, with an upper bound of $\sum m_{\nu}<0.096$ eV at 95% C.L. We find that the JBP model is consistent with $\Lambda$CDM within $2\sigma$ while the BA model sits near the outer limit of $2\sigma$. Both models exhibit a shift toward $\Lambda$CDM when the neutrino mass is included in the analysis as a free parameter. This is significant, as it weakens the apparent preference for DDE models seen in analyses where the total neutrino mass is fixed at $\sum m_{\nu}=0.06$ eV.
By mapping the thawing model into the $w_{0}$$w_{a}$ plane, we also demonstrate that this class of quintessence models is fully compatible with the regions occupied by the BA and JBP models, even extending into the phantom regime, which appears to be favored by the data. This suggests that, contrary to the claims of the DESI collaboration of phantom dark energy evidence, the data may in fact be suggesting evidence for quintessence scenarios. Notably, the thawing model also shows a moderate Bayesian preference over the CPL, with $\Delta\ln{B}=2.66$.
Finally, we studied the redshift evolution of the dark energy equation of state. Although they all remain in the quintessence regime at low redshifts, their behaviors diverge at higher redshifts. The thawing model stands out because of its smooth transition from a cosmological constant-like state ($w=-1$) at high redshift to a quintessence-like behavior near the present epoch. Overall, our results provide evidence that thawing quintessence models are not only fully compatible with widely used phenomenological parameterizations when projected in the same parameter space, but also offer tighter bounds on the neutrino mass. The thawing model also shows a strong preference for the normal mass ordering and for positive, nonzero values of $\sum m_{\nu}$.